OpenStudy (anonymous):

Compute the limit: (a) \[L=\lim_{x \rightarrow +\infty}(1+\frac{ 6 }{ x })^{x/9}\] and (b)\[L=\lim_{x \rightarrow +\infty}(1+\frac{ x }{ 7 })^{\frac{ 2 }{ \ln(x) }}\]

OpenStudy (irishboy123):

i'd bomb straight in with Bernoulli's discovery - you know the e thing so if \(\lim\limits_{n \to \infty} (1 + \dfrac{1}{n})^ n = e\) check out that on wiki then use some subs

OpenStudy (irishboy123):

like \(\dfrac{1}{n} = \dfrac{ 6 }{ x }\)

OpenStudy (irishboy123):

for the second one, instinct suggests working out the limit of the log of that function first

OpenStudy (loser66):

Still need help?

OpenStudy (anonymous):

you got this?

OpenStudy (anonymous):

no....

OpenStudy (anonymous):

\[\left(1+\frac{6}{x}\right)^{\frac{x}{9}}\]is, by definition \[\huge e^{\frac{x}{9}\log(1+\frac{6}{x})}\]

OpenStudy (anonymous):

find \[\huge \lim_{x\to \infty}\frac{x}{9}\log(1+\frac{6}{x})\] by lhopital

OpenStudy (anonymous):

then your answer will be \(e\) to that power

OpenStudy (anonymous):

do you know how to find that limit?

OpenStudy (anonymous):

I'm struggling to find the limit..

OpenStudy (anonymous):

ok lets do it and first off, lets ignore the 9 in the denominator, it is just a number, and we can put in in last

OpenStudy (anonymous):

so now the question is \[\lim_{x\to \infty}x\log(1+\frac{6}{x})\]

OpenStudy (anonymous):

the "form" is \[\infty\times 0\] is that clear or no?

OpenStudy (anonymous):

not really..

OpenStudy (anonymous):

ok so lets see that it is the form\[\lim_{x\to \infty}=\infty\] right?

OpenStudy (anonymous):

damn typo, i meant \[\lim_{x\to \infty}\color{red}x=\infty\]

OpenStudy (anonymous):

okay, I got that

OpenStudy (anonymous):

and \[\lim_{x\to \infty}\frac{6}{x}=0\]right?

OpenStudy (anonymous):

Yes. Oh yeah, 1/infinity =0. I remember that now

OpenStudy (anonymous):

making \[\lim_{x\to\infty}1+\frac{6}{x}=1\]

OpenStudy (anonymous):

and \[\log(1)=0\]so \[\lim_{x\to \infty}\log(1+\frac{6}{x})=\log(1)=0\]

OpenStudy (anonymous):

so is it now more or less clear that the form is \[\infty\times 0\]?

OpenStudy (anonymous):

okay i got that part now

OpenStudy (anonymous):

ok next, what to do with that? we can't use lhopital at the moment because it is not \[\frac{0}{0}\] or \[\frac{\infty}{\infty}\]so we do a trick of algebra

OpenStudy (anonymous):

rewrite \[x\log(1+\frac{6}{x})\]as \[\frac{\log(1+\frac{6}{x})}{\frac{1}{x}}\]

OpenStudy (anonymous):

is that algebra clear? make sure you get it, it is a common gimmick

OpenStudy (anonymous):

I got that. Oh, we're creating a ratio so that we could use the l'hopital rule

OpenStudy (anonymous):

exactly, and now it is in the form \[\frac{0}{0}\] take the derivative top and bottom separately. don't simplify too much at this step

OpenStudy (anonymous):

you got that or no?

OpenStudy (anonymous):

is it this? \[(\frac{ -6 }{ x(1+\frac{ 6 }{ x }) })+\ln(1+\frac{ 6 }{ x })/(\frac{ 1 }{ x })^{2}\]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

the derivative of \(\frac{1}{x}\)is \[-\frac{1}{x^2}\]

OpenStudy (anonymous):

the derivative of the log of something is , by the chain rule, one over something, times the derivative of something

OpenStudy (anonymous):

so the derivative of the numerator is \[\frac{1}{1+\frac{6}{x}}\times -\frac{6}{x^2}\]

OpenStudy (anonymous):

the \(-x^2\) terms cancel, leaving only \[\frac{6}{1+\frac{6}{x}}\]

OpenStudy (anonymous):

now it should be pretty clear, after you figure out that algebra, that \[\lim_{x\to \infty}\frac{6}{1+\frac{6}{x}}=6\]

OpenStudy (anonymous):

don't forget at this step that we still have to divide by that 9 we ignored out front, so your actual limit is \[\frac{6}{9}=\frac{2}{3}\]

OpenStudy (anonymous):

okay... Still kinda confused.. Is it possible to use log differentiation for this limit?

OpenStudy (anonymous):

I got this... \[1/81\ln(1+\frac{ 6 }{ x })+(\frac{ -2 }{ 3x(1+\frac{ 6 }{ x }) })\]

OpenStudy (anonymous):

where does the + sign come from?

OpenStudy (anonymous):

the product rule?

OpenStudy (anonymous):

\[\frac{d}{dx}\log(f(x))=\frac{f'(x)}{f(x)}\]

OpenStudy (anonymous):

there is no product in the numerators, it is a composition of functions \[\log(1+\frac{6}{x})\]

OpenStudy (anonymous):

the derivative is therefore \[\frac{1}{1+\frac{6}{x}}\times \frac{d}{dx}(1+\frac{6}{x})\] \[=\frac{1}{1+\frac{6}{x}}\times( -\frac{6}{x^2})\]

OpenStudy (anonymous):

Can we do this? \[\lim_{x \rightarrow +\infty}lny=\frac{ x }{ 9 }\ln(1+\frac{ 6 }{ x })\]

OpenStudy (anonymous):

yes, that is exactly what we are doing

OpenStudy (anonymous):

oh... okay.. i got this now... \[\frac{ \ln(1+\frac{ 6 }{ x }) }{ 9 }-\frac{ 2 }{ 3x(1+\frac{ 6 }{ x }) }\]

OpenStudy (anonymous):

you still have to use lhopitals rule as written above to compute that limit it looks like you took the derivative the whole thing

OpenStudy (anonymous):

all the work is written above, take a look and see if you can understand the steps none were skipped

OpenStudy (anonymous):

I am confused... :(